# Mechanistically Interpreting a Transformer-based 2-SAT Solver: An Axiomatic Approach

### v1

@@ 4 Data and Model Details
The 2-SAT Problem. Given a Boolean formula, i.e., a conjunction of disjunctive clauses over exactly two literals, the 2-SAT problem is to determine whether there is an assignment to the formula’s variables that makes the formula evaluate to true; the formula is said to be satisfiable, or SAT. For example, formula $(x_{0}\vee x_{1})\wedge(x_{1}\vee\neg x_{2})$ is satisfiable, with a satisfying assignment $x_{0}=true,x_{1}=false,x_{2}=false$ (written also as $x_{0},\neg x_{1},\neg x_{2}$). If a formula has no satisfying assignment, it is said to be unsatisfiable, or UNSAT.
@@
Dataset. We construct a dataset of randomly generated 2-SAT formulas with ten clauses and up to five variables, eliminating syntactic duplicates. We use a solver (Z3 [19]) to check satisfiability for each formula. We built a balanced dataset of $10^{6}$ SAT and $10^{6}$ UNSAT instances; we used 60% (also balanced) for training, and the rest for testing. For model analysis, we construct a separate dataset with $10^{5}$ SAT and $10^{5}$ UNSAT instances split as before in to train and test sets.
@@
5 Mechanistic Interpretability Analysis
@@
The network can be naturally decomposed into its layers and we describe the analysis of each in the following two sections. As our goal is to understand the algorithm used by the model to determine satisfiability, we refrain from analyzing behavior which does not affect the final prediction. Thus, we focus on the model’s prediction at the readout position.
5.1 First Layer as Parser
@@
Figure 1: Attention scores for all heads on the first sample of the test set. x-axis represents the source (key) positions and y-axis represents the destination (query) positions for the attention mechanism.
@@
To reverse-engineer the first layer’s behavior, we start by examining the attention patterns, i.e., the attention scores (by default, post-softmax unless stated otherwise) calculated by both layers on given input formulas. Figure 1 shows the attention scores on a formula from our test set. We clearly see patterns emerge—most tokens in the first layer attend heavily to the first token, while tokens in position $24i+$ 2, $100\leq i<$ 10, that we refer to as the second literal positions, primarily attend to themselves and the previous token, i.e., to the two literals in each clause. [^3] Recall that our formulas are of the form, $(x_{0}x_{1})(x_{1}\neg x_{2}):s$, so token positions $14i+$ 1 and $24i+$ 2 correspond to the literals of the $i^{\text{th}}$ clause in the formula. Attention from the readout token is nearly uniform. This suggests that the first layer parses each clause, storing the parsed clause information in the embeddings computed for the second literal positions.
Figure 2: Average attention scores, grouped by source token position, for all heads, calculated over the test set. In the first layer, we further average across destination token positions restricted to positions $24i+$ 2, where $100\leq i<$ 10 is the clause index. For the second layer, we only consider the readout token as the destination.
@@
In the second layer, the attentions heavily focus on these second literal positions, suggesting that the model uses the parsed clause information contained in these embeddings to classify. This allows us to restrict our analysis of the first layer to the second literal positions and the readout position. To further validate this decision, we compute the average attention score from each head on each token across the test set (see Figure 2). The results show strong periodicity with period 4, following the pattern seen in Figure 1.
Figure 3: Expected attention scores by position for the second token of each clause.
Distributional Attention Analysis. To further understand the parsing behavior noticed in Figure 1, we perform a more detailed analysis: instead of analyzing the attention patterns on individual formulas, we compute the attention scores in expectation assuming a uniform distribution over literals at each position $14i+$ 1 and $104i+2,0\leq i<$ 10. We calculate the expected pre-softmax attention scores and then apply softmax. The calculation of these expected scores is described in Appendix B.1 and the expected attention scores are shown in Figure 3, which provides further evidence of the parsing behavior seen earlier in Figures 1 and 2. A decomposed analysis of the attention scores that accounts for positional factors as well a worst-case analysis of the attention scores that consider the minimum possible attention score assigned to embeddings in position $14i+$ 1 and $24i+$ 2 by the embedding in position $24i+$ 2 further confirms the parsing hypothesis for the first layer (presented in Appendix B.1).
Interpretation. Based on our analysis of the attention patterns for the first layer, we hypothesize that it is parsing the input sequence of tokens into a list of clauses (Listing LABEL:fig:layer1_code in Appendix B.1). Note that we make this hypothesis without analyzing the MLP in the first layer. We show in Section 5.1.1 that we can successfully validate this hypothesis using our axioms. We get lucky here with the first layer—attention patterns turn out to be a less effective tool for analyzing the second layer and we are compelled to analyze the much harder-to-interpret MLP computation.
@@ 5.1.1 Checking the Axioms
Since the input to the neural model is a sequence of token IDs, $\alpha_{0}$ here is the function that translates token IDs in to tokens. To derive $\alpha_{1}$ and $\gamma_{1}$, we first compute a representative embedding for each possible clause $\psi$. Drawing from the observation that the output of attention on the second token of each clause is largely a function of the two tokens in the clause alone (ignoring all other past tokens), we compute the embedding output by layer 1 for clause $\psi$ in each position $i$ by masking attention such that it is only applied to the left and right literals in the clause. The final representative embedding for $\psi$ is the average of its embeddings in each position $i$. $\alpha_{1}$ maps the list of embeddings output by the first layer to a list of clauses by comparing (via cosine similarity) the embeddings in the second-variable positions $24i+$ 2 with the representative clause embeddings and returning the clauses with the most-similar representative embeddings. $\gamma_{1}$ maps a list of clauses to a list of embeddings. For positions other than $24i+$ 2, $\gamma_{1}$ returns the mean embedding in that position across the training set. For position $24i+$ 2, it returns the representative embedding for the clause in position $i$.
@@
Prefix Equivalence. For each clause position $i$, $\alpha_{1}$ outputs a clause $\phi:=l\lor r$, where $l$ and $r$ are literals, and, as such, we consider the clauses $r\lor l$ and $l\lor r$ equal. With this notion of equality, we obtain an $\epsilon$ [^4] We always report $\epsilon$ in terms of the upper bound of a one-sided 95% Clopper-Pearson confidence interval [13]. of approximately 0.0000374 (corresponding to perfect matching on the test set) between the abstract states output by $\alpha_{1}\circ d_{t}[1]$ and $d_{h}[1]\circ\alpha_{0}$. When we enforce order consistency, we obtain a much worse $\epsilon$ of 0.849; however, as the abstract state is order independent and the second layer’s behavior does not vary significantly with the order of variables in a clause, an order-dependent notion of equality is not necessary here.
@@
Component Equivalence. As this is the first layer, component equivalence and prefix equivalence are identical, hence the results are the same as above.
@@
Prefix Replaceability. We obtain $\epsilon\approx 0.0418$, i.e. substituting the first layer of the model with $\gamma_{1}\circ d_{h}[1]\circ\alpha_{0}$ affects the final output 4.2% of the time. In particular, this quantifies the effect of replacing the actual embedding of ‘:’ with the mean embedding across the dataset and the effect of ignoring attention to previous clauses when computing the representative clause embeddings.
@@
Component Replaceability. As this is the first layer, component and prefix replaceability are identical, hence the results are the same as above.
5.2 Second Layer as Evaluator
From the first layer’s analysis, we know that the inputs to the second layer can be described, in an abstract sense, as the list of clauses in the formula. As the model’s output is a label SAT or UNSAT, the interpretation of the second layer is a function that checks satisfiability given a list of clauses. This means that the core logic of the model is encoded primarily in the second layer.
Identifying the Key Pathway. We observe that the unembeding vector for UNSAT is nearly the negative of that of SAT, and hence, their logits are likewise negatives of each other. Also, for all formulas in the dataset, either the SAT or the UNSAT logit output at the readout position is much larger than the logits for all other tokens. Hence, the effect of components on classification can be fully described by their effects on the SAT logit. Furthermore, we observe that the effect of the attention mechanism on the SAT logit is consistently supressive and varies little between samples. Hence, the core pathway necessary to identify SAT flows through the MLP at the readout position. Unlike the first layer, we cannot identify an interpretation directly from the behavior of attention. See Appendix B.2 for more details.
Figure 4: Output coefficients of hidden neurons computed using the composition of the weight matrix for the output layer of the MLP and the unembedding matrix projected to the SAT logit.
Sparsity in the MLP Hidden Layer. We begin our analysis of the MLP by noting that, somewhat surprisingly, it exhibits sparsity in the hidden neurons. We conclude this by first observing that the second layer of the MLP and the unembedding operation are both linear, so their composition is again linear. From this perspective, then, the MLP’s effect on the classification is fully captured by a weighted sum of the post-activation outputs of the hidden neurons. We refer to these weights as neuron output coefficients for the SAT logit. This is the sense in which we observe sparsity; to see this clearly, see Figure 5. There are only 35 neurons which have a non-negligible absolute coefficient ($>10^{-6}$); one of these has a significantly lower coefficient than the others along with consistently low activation values on the samples in our analysis training set resulting in a negligble effect on the SAT logit. Hence, we restrict our analysis to the remaining 34 neurons.
Evaluation in the Hidden Neurons. Note furthermore that all the non-negligible coefficients are positive; hence, the effect of each of the relevant neurons is purely to promote SAT. This suggests that the neurons may each recognize some subset of SAT formulas instead of recognizing characteristics of UNSAT formulas. Figure 5 shows that the typical activation of one such neuron on SAT formulas is highly dependent on the assignment to the variables which satisfies the formula. This suggests that the MLP may implement a very simple algorithm: that of exhaustive evaluation. In particular, if we treat the formula $\phi$ as a Boolean function $f_{\phi}$, $\phi$ is SAT if and only if $f_{\phi}$ is not uniquely false. Hence, we can identify SAT formulas by a brute-force technique which computes the truth table of the corresponding Boolean functions and checks whether any entry is True.
A natural hypothesis is then that each of the evaluating neurons evaluates $\phi$ on some set of assignments, with some overlap between neurons; we refer to such interpretations of neuron activation behavior as disjunction-only since they take the form $h::=\phi[...]~{}|~{}h\vee h$. However, we observe that the actual behavior is less intuitive: in some cases, satisfiability with an assignment can decrease activation on some evaluating neurons. Hence, we also consider more general Boolean expressions in terms of the ϕ[]italic-$\phi[...]$’s as neuron interpretations. We find that our axiomatic analysis provides a clear way to quantify the tradeoff between interpretability and fidelity of the different interpretations.
Finally, we observe that the output coefficients of the evaluating neurons for the SAT logit are consistently high (> 2.9 in all cases), which means that a high activation on any individual evaluating neuron forces prediction of SAT; see Appendix B.2 for more details. In this sense, the action of MLP’s output layer and the unembedding layer is conceptually close to logical OR operation. We now have enough information to describe our interpretation of the second layer.
Interpretation. Based on our analysis, we decompose the interpretation of the second layer in to two components. We hypothesize that the second layer up to the MLP’s hidden layer evaluates the input formula with different assignments and outputs the results of these evaluations as a list of Booleans. The neuron interpretations describe the precise combination of assignments considered for computing each Boolean output (see Tables 1 and 2 of Appendix C for two different neuron interpretations). We evaluated the completeness of the neuron interpretations, i.e., their ability to correctly evaluate all 2-SAT formulas, using Z3. The rest of the model, namely, the MLP’s output layer and the unembedding layer, performs a logical OR over the Booleans output by the previous component and outputs True if the formula is SAT. Listing LABEL:fig:layer2_code and LABEL:fig:abstract_model in Appendix B.2 describe the mechanistic interpretation of the second layer and the entire model, respectively, using Python syntax.
5.2.1 Checking the Axioms: Hidden Layer
@@
The concrete state at the MLP’s hidden layer is a pair consisting of the residual from attention and the post-activation outputs of the hidden neurons; $\alpha_{2}$ maps these to a Boolean vector representing whether or not sufficiently high activation (> 0.5) has occurred at each of the 34 evaluating neurons. $\gamma_{2}$ simply returns a constant value for the residual (the mean on the analysis training set); the MLP’s hidden neurons have zero activation except for the evaluating neurons with True values in the corresponding position in the abstract state vector, which we assign an arbitrary large activation (2 in our experiments). Note that this is higher than the threshold for $\alpha_{2}$; we observe that this amplification is necessary for a reliable interpretation; see below.
@@
Prefix Equivalence. We obtain $\epsilon\approx 0.182$ when using the decision tree neuron interpretations, compared with approximately 0.309 for the disjunction-only interpretations. This means that the disjunction-only interpretations are much worse at predicting the intermediate state of the model. This is the only place where this simplifying choice has a cost with respect to the validity of the mechanistic interpretation; aside from component equivalence below, the disjunction-only interpretations differ little from the decision trees in all other respects. This illustrates the need to check the axiomatic properties at component level for the model under analysis.
@@
Component Equivalence. We again obtain $\epsilon\approx 0.182$ for the decision tree interpretations and 0.309 for the disjunction-only interpretations. The similarity in the $\epsilon$ values for prefix and component equivalence is because the interpretation of the first layer perfectly matches the behavior of neural network, up to literal order, on the analysis test set.
@@
Prefix Replaceability. In both cases, substituting the first two concrete layers with the first two abstract layers has minimal effect on the model’s output behavior: specifically, we get $\epsilon\approx 0.0128$ for the decision tree interpretations, and $\epsilon\approx 0.00290$ for the disjunction-only interpretations. If we drop the amplification step discussed above for $\gamma_{2}$, we significantly affect the model’s predictions: we get an $\epsilon$ of approximately 0.249 (i.e. over 24% of samples are affected) for the decision tree interpretations and approximately 0.135 for the disjunction-only interpretations.
@@
5.2.2 Checking the Axioms: Output Layer
For our analysis, we consider the model to be composed with a function which returns True if and only if the top logit at the readout position is SAT. The output type of this composed model and the mechanistic interpretation are identical; hence $\alpha_{3}$ and $\gamma_{3}$ are both the identity function.
@@ 5.2.2 Checking the Axioms: Output Layer
Prefix Replaceability. As this is the output layer of the network and $\alpha_{3}$ and $\gamma_{3}$ are the identity, this is the same as prefix equivalence.

### v2

@@ 4.1 Data and Model Details
The 2-SAT Problem. Given a Boolean formula, i.e., a conjunction of disjunctive clauses over exactly two literals, the 2-SAT problem is to determine whether there is an assignment to the formula’s variables that makes the formula evaluate to True; the formula is said to be satisfiable, or SAT. For example, the formula $(x_{0}\vee x_{1})\wedge(x_{1}\vee\neg x_{2})$ is satisfiable, with a satisfying assignment $x_{0}=\text{True},x_{1}=\text{False},x_{2}=\text{False}$ (written also as $x_{0},\neg x_{1},\neg x_{2}$). If a formula has no satisfying assignment, it is said to be unsatisfiable, or UNSAT.
@@
Dataset. We construct a dataset of randomly generated 2-SAT formulas with ten clauses and up to five variables, eliminating syntactic duplicates. We use a solver (Z3 (De Moura & Bjørner, 2008)) to check satisfiability for each formula. We built a balanced dataset of $10^{6}$ SAT and $10^{6}$ UNSAT instances; we used 60% (also balanced) for training, and the rest for testing. For model analysis, we construct a separate dataset with $10^{5}$ SAT and $10^{5}$ UNSAT instances split as before in to train and test sets.
@@
4.2 Mechanistic Interpretability Analysis
@@
The network can be naturally decomposed into its blocks and we describe the analysis of each in the following two sections. We note that our axioms help guide our analysis, particularly for the second block. As our goal is to understand the algorithm used by the model to determine satisfiability, we refrain from analyzing behavior which does not affect the final prediction. Thus, we focus on the model’s prediction at the readout position. Figure 5 in Appendix D describes our decomposition of the model into the concrete components $d_{t}[i]$.
4.2.1 First Block as Parser
@@
Figure 1: Attention scores for all heads on the first sample of the test set, the formula ($(\neg x_{0}\neg x_{1})(x_{1}\neg x_{4})(x_{1}x_{2})(x_{0}x_{3})(\neg x_{2}\neg x_{3})(x_{2}\neg x_{4})(\neg x_{0}\neg x_{3})(x_{0}x_{2})(x_{1}\neg x_{2})(x_{1}x_{4})$. The x-axis represents the source (key) positions and the y-axis represents the destination (query) positions for the attention mechanism.
@@
To reverse-engineer the first block’s behavior, we start by examining the attention patterns, i.e., the attention scores (by default, post-softmax unless stated otherwise) calculated by both blocks on given input formulas. Figure 1 shows the attention scores on a formula from our test set. We clearly see patterns emerge—most tokens in the first block attend heavily to the first token, while tokens in position $24i+$ 2, $100\leq i<$ 10, that we refer to as the second literal positions, primarily attend to themselves and the previous token, i.e., to the two literals in each clause. [^3] Formulas are of the form, $(x_{0}x_{1})(x_{1}\neg x_{2}):s$, so token positions $14i+$ 1 and $24i+$ 2 correspond to the literals of the $i^{\text{th}}$ clause. Attention from the readout token is nearly uniform. This suggests that the first block parses each clause, storing the parsed clause information in the token representations output by the first block in the second literal positions.
@@
In the second block, the attentions heavily focus on these second literal positions, suggesting that the model uses the parsed clause information contained in these representations to classify. This allows us to restrict our analysis of the first block to the second literal positions and the readout position. To further validate this decision, we compute the average attention score from each head on each token across the test set (see Figure 6 in Appendix D.1). The results show strong periodicity with period 4, following the pattern seen in Figure 1.
We perform additional distributional and worst-case analysis of attention patterns (described in Appendix D.1) and these analyses confirm the parsing hypothesis for the first block.
Interpretation. Based on our analysis of the attention patterns for the first block, we hypothesize that it is parsing the input sequence of tokens into a list of clauses (Listing LABEL:fig:layer1_code in Appendix D.1). The behavior of attention is enough to form this hypothesis, as the sparse attention patterns show that information primarily flows from tokens in a clause to the token in the second literal position in the same clause; hence we can view the action of the first block as computing a representation for each clause in the input. In this sense, analyzing the MLP in the first block is not necessary to identify the semantic function of the first block.
We show in Sec 4.2.3 that we can successfully validate this hypothesis using our axioms. As we’ll see in the analysis of the second block in Sec 4.2.2, we cannot always count on self-explanatory attention patterns; for the second block, we will be forced to analyze the much harder-to-interpret MLP.
4.2.2 Second Block as Evaluator
From the first block’s analysis, we know that the inputs to the second block can be described, in an abstract sense, as the list of clauses in the formula. As the model’s output is a label SAT or UNSAT, the interpretation of the second block is a function that checks satisfiability given a list of clauses. This means that the core logic of the model is encoded primarily in the second block.
Identifying the Key Pathway. We observe that the unembedding vector for UNSAT is nearly the negative of that of SAT (in particular, $\|W_{U}^{\text{SAT}}+W_{U}^{\text{UNSAT}}\|<10^{-5}$ and their cosine similarity is -1 up to floating point precision), and hence, their logits are likewise negatives of each other. Also, for all formulas in the dataset, either the SAT or the UNSAT logit output at the readout position is much larger than the logits for all other tokens. Hence, the effect of components on classification can be fully described by their effects on the SAT logit. Furthermore, we observe that the effect of the attention mechanism on the SAT logit is consistently suppressive and varies little between samples. Hence, the core pathway necessary to identify SAT flows through the MLP at the readout position. Unlike the first block, we cannot identify an interpretation directly from the behavior of attention. See Appendix D.2 for more details.
Figure 2: Output coefficients of hidden neurons computed via the composition of the output layer of the MLP and the unembedding matrix projected to the SAT logit.
Sparsity in the MLP Hidden Layer. We begin our analysis of the MLP by noting that, somewhat surprisingly, it exhibits sparsity in the hidden neurons. We conclude this by first observing that the second layer of the MLP and the unembedding operation are both linear, so their composition is again linear. From this perspective, then, the MLP’s effect on the classification is fully captured by a weighted sum of the post-activation outputs of the hidden neurons. We refer to these weights as neuron output coefficients for the SAT logit. This is the sense in which we observe sparsity; to see this clearly, see Figure 2. There are only 35 neurons which have a non-negligible absolute coefficient ($>10^{-6}$); one of these has a significantly lower coefficient than the others along with consistently low activation values on the samples in our analysis training set resulting in a negligible effect on the SAT logit. Hence, we restrict our analysis to the remaining 34 neurons.
Evaluation in the Hidden Neurons. Note furthermore that all the non-negligible coefficients are positive; hence, the effect of each of the relevant neurons is purely to promote SAT. This suggests that the neurons may each recognize some subset of SAT formulas instead of recognizing characteristics of UNSAT formulas. Figure 3 shows that the typical activation of one such neuron on SAT formulas is highly dependent on the assignment to the variables which satisfies the formula. This suggests that the MLP may implement a very simple algorithm: that of exhaustive evaluation. In particular, if we treat the formula $\phi$ as a Boolean function $f_{\phi}$, $\phi$ is SAT if and only if $f_{\phi}$ is not the constant False function. Hence, we can identify SAT formulas by a brute-force technique which computes the truth table of the corresponding Boolean functions and checks whether any entry is True.
A natural hypothesis is then that each of the evaluating neurons evaluates $\phi$ on some set of assignments, with some overlap between neurons; we refer to such interpretations of neuron activation behavior as disjunction-only since they take the form $h::=\phi[...]~{}|~{}h\vee h$. However, we observe that the actual behavior is less intuitive: in some cases, satisfiability with an assignment can decrease activation on some evaluating neurons. Hence, we also consider more general Boolean expressions in terms of the ϕ[]italic-$\phi[...]$’s as neuron interpretations. We find that our axiomatic analysis provides a clear way to quantify the trade-off between interpretability and fidelity of the different interpretations.
Finally, we observe that the output coefficients of the evaluating neurons for the SAT logit are consistently high (> 2.9 in all cases), which means that a high activation on any individual evaluating neuron forces prediction of SAT; see Appendix D.2 for more details. In this sense, the action of MLP’s output layer and the unembedding layer is conceptually close to logical OR operation. We now have enough information to describe our interpretation of the second block.
Interpretation. Based on our analysis, we decompose the interpretation of the second block into two components. We hypothesize that the second block up to the MLP’s hidden layer evaluates the input formula with different assignments and outputs the results of these evaluations as a list of Booleans. The neuron interpretations describe the precise combination of assignments considered for computing each Boolean output (see Tables 2 and 3 of Appendix G for two different neuron interpretations). We evaluated the completeness of the neuron interpretations, i.e., their ability to correctly evaluate all 2-SAT formulas, using Z3 (De Moura & Bjørner, 2008). The rest of the model, namely, the MLP’s output layer and the unembedding layer, performs a logical OR over the Booleans output by the previous component and outputs True if the formula is SAT. Listing LABEL:fig:layer2_code and LABEL:fig:abstract_model in Appendix D.2 describe the mechanistic interpretation of the second block and the entire model, respectively, using Python syntax.
@@ 4.2.3 Checking the Axioms
First Block. Since the input to the neural model is a sequence of token IDs, $\alpha_{0}$ here is the function that translates token IDs into tokens. To derive $\alpha_{1}$ and $\gamma_{1}$, we first compute a canonical representation for each possible clause $\psi$.
@@
Drawing from the observation that the output of attention on the second token of each clause is largely a function of the two tokens in the clause alone (ignoring all other past tokens), we compute the representation output by block 1 for clause $\psi$ in each position $i$ by masking attention such that it is only applied to the left and right literals in the clause. The canonical representation for $\psi$ is the average of the representations derived for each position $i$. $\alpha_{1}:[\texttt{Tensor}]\rightarrow[(\texttt{Literal},\texttt{Literal})]$, where Literal is the type consisting of the $x_{i}$ and their negations, maps the list of representations output by the first block to a list of clauses by comparing (via cosine similarity) the representations in the second-variable positions $24i+$ 2 with the canonical clause representations and returning the clauses with the most-similar representations. $\gamma_{1}:[(\texttt{Literal},\texttt{Literal})]\rightarrow[\texttt{Tensor}]$ maps a list of clauses to a list of canonical clause representations. For positions other than $24i+$ 2, $\gamma_{1}$ returns the mean representation in that position across the training set. For position $24i+$ 2, it returns the canonical representation for the clause in position $i$. See Listing LABEL:lst:alphas_1 and Listing LABEL:lst:gammas_1 of Appendix D for pseudocode for $\alpha_{1}$ and $\gamma_{1}$, respectively.
@@
Prefix Equivalence. For each clause position $i$, $\alpha_{1}$ outputs a clause $\phi:=l\lor r$, where $l$ and $r$ are literals, and, as such, we consider the clauses $r\lor l$ and $l\lor r$ equal. With this notion of equality, we obtain an $\epsilon$ [^4] We always report $\epsilon$ in terms of the upper bound of a one-sided 95% Clopper-Pearson confidence interval. of approximately 0.0000374 (corresponding to perfect matching on the test set) between the abstract states output by $\alpha_{1}\circ d_{t}[1]$ and $d_{h}[1]\circ\alpha_{0}$. When we enforce order consistency, we obtain a much worse $\epsilon$ of 0.849; however, as the abstract state is order independent and the second block’s behavior does not vary significantly with the order of variables in a clause, an order-dependent notion of equality is not necessary here.
@@
Component Equivalence. As this is the first component, component equivalence and prefix equivalence are identical, hence the results are the same as above.
@@
Prefix Replaceability. We obtain $\epsilon\approx 0.0418$, i.e. substituting the first component of the model with $\gamma_{1}\circ d_{h}[1]\circ\alpha_{0}$ affects the final output 4.2% of the time. In particular, this quantifies the effect of replacing the actual first-block representation of ‘:’ with its mean value across the dataset and the effect of ignoring attention to previous clauses when computing the canonical clause representations.
@@
Component Replaceability. As this is the first component, component and prefix replaceability are identical, hence the results are the same as above.
@@
Hidden Layer of Second Block. The output of the second concrete component is a pair consisting of the residual from attention and the post-activation outputs of the hidden neurons; $\alpha_{2}:(\texttt{Tensor},\texttt{Tensor})\rightarrow[\texttt{bool}]$ maps these to a Boolean vector representing whether or not sufficiently high activation (> 0.5) has occurred at each of the 34 evaluating neurons. $\gamma_{2}:[\texttt{bool}]\rightarrow(\texttt{Tensor},\texttt{Tensor})$ simply returns a constant value for the residual (the mean on the analysis training set); the MLP’s hidden neurons have zero activation except for the evaluating neurons with True values in the corresponding position in the abstract state vector, which we assign an arbitrary large activation (2 in our experiments). Note that this is higher than the threshold for $\alpha_{2}$; we observe that this amplification is necessary for a reliable interpretation; see below. See Listing LABEL:lst:alphas_2 and Listing LABEL:lst:gammas_2 of Appendix D for pseudocode for $\alpha_{2}$ and $\gamma_{2}$, respectively.
@@
Prefix Equivalence. We obtain $\epsilon\approx 0.182$ when using the decision tree neuron interpretations, compared with approximately 0.309 for the disjunction-only interpretations in which our neuron models activate on formulas satisfiable with any of a set of assignments. This means that the disjunction-only interpretations are much worse at predicting the intermediate state of the model. This is the only place where this simplifying choice has a cost with respect to the validity of the mechanistic interpretation; aside from component equivalence below, the disjunction-only interpretations differ little from the decision trees in all other respects. This illustrates the need to check the axiomatic properties at component level for the model under analysis.
@@
Component Equivalence. We again obtain $\epsilon\approx 0.182$ for the decision tree interpretations and 0.309 for the disjunction-only interpretations. The similarity in the $\epsilon$ values for prefix and component equivalence is because the interpretation of the first component perfectly matches the behavior of neural network, up to literal order, on the analysis test set.
@@
Prefix Replaceability. In both cases, substituting the first two concrete components with the first two abstract components has minimal effect on the model’s output behavior: specifically, we get $\epsilon\approx 0.0128$ for the decision tree interpretations, and $\epsilon\approx 0.00290$ for the disjunction-only interpretations. If we drop the amplification step discussed above for $\gamma_{2}$, we significantly affect the model’s predictions: we get an $\epsilon$ of approximately 0.249 (i.e. over 24% of samples are affected) for the decision tree interpretations and approximately 0.135 for the disjunction-only interpretations.
@@
Output Layer of Second Block. For our analysis, we consider the model to be composed with a function which returns True if and only if the top logit at the readout position is SAT. The output type of this composed model and the mechanistic interpretation are identical; hence $\alpha_{3}$ and $\gamma_{3}$ are both the identity function.
@@
Prefix Replaceability. As this is the final component of the network and $\alpha_{3}$ and $\gamma_{3}$ are the identity, this is the same as prefix equivalence.
@@
5 Case Study: Modular Addition
Nanda et al. (2023a) train a model with a single Transformer block to perform modular addition. The model is given input of the form $a\ b\ =$ and returns $a+b\mod P$ at the ‘=’ token where $P$ is fixed to be 113. They claim the following mechanistic interpretation of the concrete model (see Listing LABEL:lst:grokking_code in Appendix H for the pseudocode for the abstract components):
  1. 1.
    Encoding of inputs: The embedding matrix represents $a$ and $b$ as $\sin(w_{k}a)$, $\sin(w_{k}b)$, $\cos(w_{k}a)$, and $\cos(w_{k}b)$ for key frequencies $w_{k}=(2k\pi)/P$ for $k\in\{14,35,41,42,52\}$.
  2. 2.
    Computation of sum-of-angles identities: The attention and MLP input layer compute $\cos(w_{k}(a+b))$ and $\sin(w_{k}(a+b))$ using trigonometric identities.
  3. 3.
    Difference-of-angles identities and argmax: The MLP output layer and the unembedding matrix computes $\cos(w_{k}(a+b-c))$ for each key frequency and for each $c\in\mathbb{Z}_{P}$, adds the computed cosines, and then selects the $c$ maximizing the sum, which occurs at $c^{*}=a+b\mod P$.
As all abstract components but the last one output continuous values, we must discretize the output of each abstract component for Axioms 1 and 2 to be meaningful. For the first component, we apply a simple discretization of rounding up to three decimal points; for the second component, we define an equivalence class which treats abstract states as equivalent when their downstream effects on both the abstract and concrete models are identical. Discretizations of this form are quite general: an equivalence class up to the next discrete-valued intermediate value is expressible for any abstract model with a discrete-valued output. Axioms 3 and 4 ensure that any discretization does not impact end-to-end behavior. The $\alpha$ and $\gamma$ functions for each component are linear maps learned from data. See Appendix H for more details.
We apply our approach to this model and confirm that the Nanda et al. (2023a)’s mechanistic interpretation satisfies our axioms, noting that, in their paper, while Nanda et al. (2023a) provide evidence similar to our Axioms 2 and 4 in support of the validity of their interpretation, they do not provide any evidence as required by our Axioms 1 and 3.
In particular, with the above discretization we obtain a very strong $\epsilon$ of 0.000335, which corresponds to perfect matching. However, it is important to note that results with simple discretizations indicate that the model’s behavior is not fully captured. In particular, we obtain an $\epsilon$ of 1 from Axioms 1 and 2 for the second component when, instead, we discretize by rounding to a single decimal place. There are important tradeoffs to the application of this type of discretization: while it enables the analyst to eliminate variation which has no impact on the downstream behavior of the model, it also limits the ability of Axioms 1 and 2 to validate the internal behavior.